# Choosing the Better Bandit Algorithm under Data Sharing: When Do A/B Experiments Work?

### v1

@@ Abstract
We study A/B experiments that are designed to compare the performance of two recommendation algorithms. Prior work has shown that the standard difference-in-means estimator is biased in estimating the global treatment effect (GTE) due to a particular form of interference between experimental units. Specifically, units under the treatment and control algorithms contribute to a shared pool of data that subsequently train both algorithms, resulting in interference between the two groups. The bias arising from this type of data sharing is known as “symbiosis bias”. In this paper, we highlight that, for decision-making purposes, the sign of the GTE often matters more than its precise magnitude when selecting the better algorithm. We formalize this insight under a multi-armed bandit framework and theoretically characterize when the sign of the expected GTE estimate under data sharing aligns with or contradicts the sign of the true GTE. Our analysis identifies the level of exploration versus exploitation as a key determinant of how symbiosis bias impacts algorithm selection.
@@ 1 Introduction
Interference refers to the phenomenon where users in the treatment and control groups influence each other, rendering the DM estimator biased for the GTE. On a production-scale recommendation system, the following field experiment has been conducted to demonstrate the effect of such interference [12]. In this experiment, the treatment to users is to recommended more content that is recently published. Accordingly, these users consume more recently-published content. However, the more interesting observation is that users in the control group, with the recommendation algorithm serving them unchanged, also start to consume more recently-published content.
This phenomenon of interference is attributed to shared data feedback loops [12], illustrated in Figure 1a (left). As users in the treatment group engage more with recently-published content, they produce training data associated with the recently-published content. Such data is shared with and used to train the control algorithm, creating a mechanism for the treatment group to influence the control group. This experiment setup thus violates SUTVA and yields biased GTE estimates. This bias arising from such shared feedback loops is termed “symbiosis bias”, describing how the two algorithms depend on each other in a symbiotic fashion through sharing data. [^1] Interactions among users across treatment and control groups may also contribute to such phenomenon. For example, users in the control group engage with social media posts that their friends in the treatment group have liked or commented on, or users in the control group make purchases because they hear about these products from their friends in the treatment group. To minimize such network effects, the experiment randomizes treatment assignment on a country level. That is, each country, along with all users in the country, is assigned to either the treatment or control groups uniformly at random. See [12, Section 5] for additional details on the precise experiment setup.
@@
Before we proceed to analyze the impact of symbiosis bias, let us consider two natural ideas to mitigate this bias and the additional challenges they introduce. The first mitigation strategy is to isolate the treatment and control groups so that they do not share any data, as shown in Figure 1a (right). This idealized design eliminates the shared feedback loops and hence the symbiosis bias, but has a number of drawbacks. First, it reduces the efficiency of learning because each algorithm now only has access to a smaller dataset for training [48], which is especially a problem if a small fraction of users is assigned to the treatment group. In business applications, it is desirable to reach decisions quickly, and using shared data to speed up learning is especially attractive in scenarios where the data from the treatment and control groups have similar distributions and the impact of symbiosis bias is minimal. Second, in large-scale recommendation systems, maintaining two separate datasets increases the engineering efforts. Using social media as an example, the content may be ranked by a scoring function that predicts the number of likes for the content. If this scoring function is trained from the number of likes on historical data, now the number of likes by users in the treatment and control groups needs to be separately counted. The scoring function may also use neural networks to extract features from images and texts, and these neural networks are fine-tuned on user data. Now two separate neural networks need to be trained and stored. Such extra computation and storage needs are not scalable as companies often run many A/B experiments in parallel.
A second mitigation strategy is to deploy the treatment or control algorithms to all users, but at different times, a design known as “switchback experiments”. However, this experiment design introduces other types of time-dependent biases [30, 9]. For example, in online marketplaces, consider a scenario where the control algorithm is deployed first and recommends a product to the user. After some time, the treatment algorithm is deployed and the user makes the purchase. It is unclear if the purchase decision is made due to the treatment algorithm or the control algorithm. Moreover, user behavior is inherently non-stationary. For instance, purchases of umbrellas increase during rainy seasons, and consumer demand in general fluctuates in response to broader economic conditions. It is thus non-trivial to design GTE estimators that carefully account for such non-stationary behavior in switchback experiments.
@@ 1.1 Overview of results
We consider multi-armed bandit algorithms as candidate recommendation algorithms. Our metric of interest for measuring the GTE is the expected cumulative regret: a higher regret indicates worse performance. We quantify the expected regret of a pair of algorithms with and without data sharing. These regret bounds, in turn, allow us to characterize the impact of symbiosis bias on the expected GTE estimate produced by the DM estimator.
@@
We first identify cases where the expected GTE estimate under data sharing has the opposite sign to the true GTE defined without data sharing, a situation we refer to as “sign violation”. Specifically, we consider the greedy algorithm which selects the arm with the highest empirical mean at each timestep. We show that while the greedy algorithm is known to be highly suboptimal and incurs a linear regret, its regret is reduced significantly to a constant under data sharing, when it runs jointly with a general class of algorithms (Theorem 3), including the classic $\epsilon$-greedy algorithm (Theorem 1) and the upper confidence bound (UCB) algorithm (Theorem 2). This result suggests that the sign of the expected GTE estimate under data sharing yields a wrong comparison of the two algorithms. Moreover, our analysis shows that the magnitude of GTE estimate under data sharing is $\Theta(\log T)$ in expectation, whereas the true GTE is $\Theta(T)$, suggesting that the DM estimator significantly under-estimates the magnitude of the GTE.
@@
We then characterize scenarios where symbiosis bias does not flip the sign of the expected GTE estimate, a regime we refer to as “sign preservation”. We consider classes of $\epsilon$-greedy and UCB algorithms parameterized by a parameter $\alpha\in[0,1]$. This parameter $\alpha$ captures the trade-off between exploration and exploitation: a higher value of $\alpha$ corresponds to more exploration. The algorithm incurs a regret of $\Theta(\log T)$ when $\alpha=0$ and a regret of $\Theta(T^{\alpha})$ when $0<\alpha\leq 1$. These parameterized classes hence cover algorithms with different performances from being optimal $\Theta(\log T)$ to the worst possible $\Theta(T)$ regret. We show that data sharing does not affect either the sign or the order of magnitude of the expected GTE estimates, for both the $\epsilon$-greedy (Theorem 4) and UCB (Theorem 5) algorithms.
@@
The intuition behind these theoretical results is illustrated in Figure 2. We characterize a bandit algorithm by its level of exploration versus exploitation. When an algorithm runs individually (left panel), too much exploration (e.g., sampling arms uniformly at random) or too much exploitation (e.g., the greedy algorithm) is highly suboptimal and leads to a linear regret. The optimal $O(\log T)$ regret is achieved by a careful balance between exploration and exploitation. In contrast, when two algorithms run jointly under data sharing (right panel), exploration becomes the “tragedy of the commons”. When an algorithm explores, the data from exploration is shared between two algorithms, but the cost to exploration (i.e., a high instantaneous regret by pulling a potentially suboptimal arm) only incurs to the algorithm that explores. Hence, algorithms have no incentive to explore, and the more it exploits, the more it takes advantage of the other algorithm. The greedy algorithm thus performs very well by free riding the exploration done by the other algorithm.
Given this intuition, the regime of sign preservation applies to a pair of algorithms that overly explore (compared to the optimal balance of exploration versus exploitation leading to a regret of $O(\log T)$), where over-exploration increases its regret with or without data sharing. In contrast, an overly-exploiting algorithm leads to sign violation, because over-exploitation (and hence insufficient exploration) is only beneficial under data sharing due to free riding, but detrimental without data sharing.
@@ 1.2.1 Causal inference
Our focus on the sign of the GTE for decision making shares similar spirit to Johari et al. [36], which studies interference due to constrained inventory in online marketplaces. Despite the fact that the difference-in-means (DM) estimator is biased for the GTE, Johari et al. [36] identify scenarios under which the estimated GTE and the true GTE have the same or opposite signs, and show that the biased DM estimator surprisingly leads to a better hypothesis test in certain cases.
@@ 1.2.2 Bandit problems
Shivaswamy and Joachims [47] consider the stochastic multi-armed bandit problem where the algorithm has access to a set of historical samples from each arm, and shows that $\Omega(\log T)$ historical samples enable the algorithm to achieve a constant regret. Variants of this problem include allowing the algorithm to receive additional samples in each iteration by paying a certain cost (where the cost is allowed to be zero) [57], or allowing the algorithm to “abstain” in an iteration, so that the the sample is observed by the algorithm but not counted towards its regret [55]. In our formulation, the two algorithms receive “auxiliary” data from each other. Contrary to prior work, the treatment algorithm is not free to choose directly what arms are pulled in the auxiliary data which is determined by the control algorithm, and vice versa, though the treatment algorithm indirectly influences the auxiliary data supplied by the control algorithm through data sharing. Such dependencies between the two algorithms do not exist when the historical data is offline.
@@ 1.2.3 Intersection between causal inference and bandit algorithms
Our work lies at the intersection of causal inference, experimentation, and bandit problems. A growing body of prior literature [59, 24, 58, 25, 40, 49] uses bandit algorithms for adaptive data collection to address causal questions that involve evaluating the performance of a policy or an arm. For evaluating policies, one objective is to perform off-policy evaluation, that is, to evaluate the policy’s performance on the true distributions, only given biased data that is collected adaptively. For evaluating arms, the objective is to estimate the mean of each arm or to construct confidence intervals for these means. This line of work uses bandit algorithms as a tool to evaluate other objects, whereas our goal is to evaluate bandit algorithms themselves. In other words, we focus on comparing adaptive algorithms rather than fixed policies or arms.
@@ 2.1 Stochastic multi-armed bandit
A bandit algorithm aims to maximize the cumulative reward over a horizon of $T\geq 1$ timesteps. At each timestep $t\in[T]$, the bandit algorithm selects an arm $I_{t}\in[K]$ and receives a stochastic reward $Y_{t}\sim D_{I_{t}}$ from this arm. The arm selection is based on the history of all arm pulls and reward samples in the previous $(t-1)$ timesteps. Formally, we denote the history at timestep $t$ by $\mathcal{H}_{t}=(\mathcal{H}_{1,t},\ldots,\mathcal{H}_{K,t})$, where $\mathcal{H}_{k,t}$ denotes the set of reward samples collected from arm $k\in[K]$ in the previous $(t-1)$ timesteps. That is, we define $\mathcal{H}_{k,t}:=\{Y_{i}:I_{i}=k,i\in[t-1]\}$. In words, we consider a class of bandit algorithms that select arms based on the set of reward samples for each arm so far, but are oblivious to the order of the past arm pulls (or reward samples). Many classic bandit algorithms belong to this general class, including the $\epsilon$-greedy algorithm, the Upper Confidence Bound (UCB) algorithm, the EXP3 algorithm, and Thompson sampling. For any algorithm $\mathcal{A}$ in this class, we write the arm selected by this algorithm at timestep $t$ by $I_{t}=\mathcal{A}(\mathcal{H}_{t})$.
@@ 2.2 Bandit algorithms under data sharing
We formulate an A/B experiment comparing two recommendation algorithms as a bandit problem with $2T$ users. We assume that two users arrive at the platform at each timestep $t\in[T]$. These two users are assigned to the treatment or the control algorithm one each uniformly at random. To model data sharing between the treatment and control algorithms, we assume that the two algorithms make recommendations based on past ratings from users in both groups. Formally, consider two bandit algorithms $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ that share history $\mathcal{H}_{t}$ over the entire course of $T$ timesteps. At each timestep $t\in[T]$, each of the two algorithms independently chooses the arm $\mathcal{A}_{1}(\mathcal{H}_{t})$ and $\mathcal{A}_{2}(\mathcal{H}_{t})$ to pull based on history $\mathcal{H}_{t}$, observes a reward sample from the chosen arm, and adds this reward sample to the shared history $\mathcal{H}_{t}$. Thus, two reward samples are added to the shared history at each timestep $t$, and the total number of samples included in history $\mathcal{H}_{t}$ is $2(t-1)$. This procedure of running two algorithms under data sharing is formalized in Algorithm 1.

### v2

@@ Abstract
We study A/B experiments that are designed to compare the performance of two recommendation algorithms. Prior work has observed that the stable unit treatment value assumption (SUTVA) often does not hold in large-scale recommendation systems, and hence the estimate for the global treatment effect (GTE) is biased. Specifically, units under the treatment and control algorithms contribute to a shared pool of data that subsequently train both algorithms, resulting in interference between the two groups. In this paper, we investigate when such interference may affect our decision making on which algorithm is better. We formalize this insight under a multi-armed bandit framework and theoretically characterize when the sign of the difference-in-means estimator of the GTE under data sharing aligns with or contradicts the sign of the true GTE. Our analysis identifies the level of exploration versus exploitation as a key determinant of how data sharing impacts decision making, and we propose a detection procedure based on ramp-up experiments to signal incorrect algorithm comparison in practice.
@@ 1 Introduction
Interference refers to the phenomenon where users in the treatment and control groups influence each other, rendering the DM estimator biased for the GTE. On a production-scale recommendation system, the following field experiment has been conducted to demonstrate the effect of such interference [13]. In this experiment, the treatment to users is to recommend more content that is recently published. Accordingly, these users consume more recently-published content. However, the more interesting observation is that users in the control group, with the recommendation algorithm serving them unchanged, also start to consume more recently-published content.
This phenomenon of interference is attributed to shared data feedback loops [13], illustrated in Figure 1 (left). As users in the treatment group engage more with recently-published content, they produce training data associated with the recently-published content. Such data is shared with and used to train the control algorithm, creating a mechanism for the treatment group to influence the control group. This experiment setup thus violates SUTVA and yields biased GTE estimates. This bias arising from such shared feedback loops is termed “symbiosis bias”, describing how the two algorithms depend on each other in a symbiotic fashion through sharing data. [^1] Interactions among users across treatment and control groups may also contribute to such phenomenon. For example, users in the control group engage with social media posts that their friends in the treatment group have liked or commented on, or users in the control group make purchases because they hear about these products from their friends in the treatment group. To minimize such network effects, the experiment randomizes treatment assignment on a country level. That is, each country, along with all users in the country, is assigned to either the treatment or control groups uniformly at random. See [13, Section 5] for additional details on the precise experiment setup.
((a))
((b))
@@
Before we proceed to analyze the impact of symbiosis bias, let us consider two natural ideas to mitigate this bias and the additional challenges they introduce. The first mitigation strategy is to isolate the treatment and control groups so that they do not share any data, as shown in Figure 1 (right). This idealized design eliminates the shared feedback loops and hence the symbiosis bias, but has a number of drawbacks. First, it reduces the efficiency of learning because each algorithm now only has access to a smaller dataset for training [58], which is especially a problem if a small fraction of users is assigned to the treatment group. In business applications, it is desirable to reach decisions quickly, and using shared data to speed up learning is especially attractive in scenarios where the data from the treatment and control groups have similar distributions and the impact of symbiosis bias is minimal. Second, in large-scale recommendation systems, maintaining two separate datasets increases the engineering efforts. Using social media as an example, the content may be ranked by a scoring function that predicts the number of likes for the content. If this scoring function is trained from the number of likes on historical data, now the number of likes by users in the treatment and control groups needs to be separately counted. The scoring function may also use neural networks to extract features from images and texts, and these neural networks are fine-tuned on user data. Now two separate neural networks need to be trained and stored. Such extra computation and storage needs are not scalable as companies often run many A/B experiments in parallel.
A second mitigation strategy is to deploy the treatment or control algorithm to all users, but at different times, a design known as “switchback experiments”. However, this experiment design introduces other types of time-dependent biases [32, 9]. For example, in online marketplaces, consider a scenario where the control algorithm is deployed first and recommends a product to the user. After some time, the treatment algorithm is deployed and the user makes the purchase. It is unclear if the purchase decision is made due to the treatment algorithm or the control algorithm. Moreover, user behavior is inherently non-stationary. For instance, purchases of umbrellas increase during rainy seasons, and consumer demand in general fluctuates in response to broader economic conditions. It is thus non-trivial to design GTE estimators that carefully account for such non-stationary behavior in switchback experiments.
@@ 1.1 Overview of results
We answer the above question through the lens of the exploration–exploitation trade-off. Our key finding is that when comparing two recommendation algorithms, the impact of symbiosis bias depends on their level of exploration: If the worse algorithm underperforms because it over-explores (without data sharing), then symbiosis bias does not alter the sign of the expected GTE estimates under data sharing (a regime that we refer to as “sign preservation”). In contrast, if the worse algorithm underperforms because it overexploits (without data sharing), then symbiosis bias reverses the sign of the expected GTE estimates (a regime that refer to as “sign violation”). We formally establish this result in terms of bandit algorithms, which play a central role in recommendation systems. Under a multi-armed bandit setup, our metric of interest for measuring the GTE is the expected cumulative regret, where a higher regret indicates worse performance.
@@
We use the greedy algorithm, which always selects the arm with the highest empirical mean, as our example of an over-exploiting bandit algorithm. While the greedy algorithm is highly suboptimal and incurs a linear regret, we show that its regret is reduced to a constant under data sharing, when it runs jointly with a general class of algorithms (Theorem 3), including the classic $\epsilon$-greedy algorithm (Theorem 1) and the upper confidence bound (UCB) algorithm (Theorem 2). This result suggests that the sign of the expected GTE estimate under data sharing yields a wrong comparison of the two algorithms. Moreover, we show that the magnitude of GTE estimate under data sharing is $\Theta(\log T)$ in expectation, whereas the true GTE is $\Theta(T)$, suggesting that the DM estimator significantly under-estimates the magnitude of the GTE.
@@
We then consider $\epsilon$-greedy and UCB algorithms as our examples of over-exploring bandit algorithms. We parameterize these algorithms by a parameter $\alpha\in[0,1]$, where a higher value of $\alpha$ corresponds to more exploration. The algorithm incurs a regret of $\Theta(\log T)$ when $\alpha=0$ and a regret of $\Theta(T^{\alpha})$ when $0<\alpha\leq 1$. These parameterized classes hence cover algorithms with different performances from being optimal $\Theta(\log T)$ to the worst possible $\Theta(T)$ regret due to over-exploration. We show that data sharing does not affect either the sign or the order of magnitude of the expected GTE estimates, for both the $\epsilon$-greedy and UCB algorithms (Theorem 4).
Beyond the greedy, $\epsilon$-greedy, and UCB algorithms, we also study EXP3 and Thompson Sampling in Appendix A and Section 5.3 respectively, where we establish similar results, either theoretically or empirically.
@@
These results together substantiate our theoretical insights on exploration versus exploitation, illustrated in Figure 2. When an algorithm runs individually (left panel), too much exploration (e.g., sampling arms uniformly at random) or too much exploitation (e.g., the greedy algorithm) is suboptimal. The optimal $O(\log T)$ regret is achieved by a careful balance between exploration and exploitation. In contrast, when two algorithms run jointly under data sharing (right panel), exploration becomes the “tragedy of the commons”. When an algorithm explores, the data from exploration is shared between two algorithms, but the cost to exploration (i.e., a high instantaneous regret by pulling a potentially suboptimal arm) only incurs to the algorithm that explores. Hence, the algorithm is incentivized to free ride the exploration done by the other algorithm rather than explore on its own.
Our results allow companies interested in selecting the better algorithm in A/B experiments to face a simpler problem: identifying whether their experiment falls under the regime of sign preservation or sign violation, rather than estimating the precise magnitude of symbiosis bias (as in, for example, [58]). In Section 4, we address the practical question of how to diagnose these regimes in real A/B experiments. Specifically, we study ramp-up experiments in which the treatment allocation probability varies over a sequence of experiments. We find that when increasing traffic to the seemingly better algorithm leads to a deterioration in overall performance, this pattern is indicative of a sign violation regime. We formalize this intuition by proposing a simple diagnostic based on the monotonicity of average outcomes and provide theoretical justification for why this criterion distinguishes sign violation from sign preservation. This result connects our theoretical insights to standard experimental practice and offers a lightweight tool for practitioners to assess whether comparisons under data sharing can be trusted.
@@ 1.2.1 Causal inference
Our focus on the sign of the GTE for decision making shares similar spirit to Johari et al. [38], which studies interference due to constrained inventory in online marketplaces. Despite the fact that the difference-in-means (DM) estimator is biased for the GTE, Johari et al. [38] identify scenarios under which the estimated GTE and the true GTE have the same or opposite signs, and show that the biased DM estimator surprisingly leads to a better hypothesis test in certain cases. Johari et al. [39] study the impact of interference on decision-making when conducting pricing experimentation. Ni [48] provides the guidance on whether to roll out or shelve the change from a robust optimization perspective.
@@
Ramp-up experiments under interference
Ramp-up experiments are widely used in industry to gradually increase exposure to new features or algorithms in order to manage risk and monitor performance, and are a standard component of large-scale experimentation platforms [64, 40, 41, 66]. Beyond their operational role, recent work has shown that ramp-up experiments can be informative in settings with interference. Ramp-up experiments are used to detect the existence of interference [26], to aid identification and model selection [12], and to estimate the global treatment effect under network interference without requiring full knowledge of the interference graph [69]. Our work develops a new use of ramp-up experiments in the context of interference induced by shared data feedback loops when comparing learning algorithms. Specifically, we use changes in overall performance during ramp-up to diagnose whether an A/B comparison operates in the regime of sign violation or sign preservation, providing guidance on when the direction of an experimental conclusion can be trusted.
@@ 1.2.2 Bandit problems
Shivaswamy and Joachims [57] consider the stochastic multi-armed bandit problem where the algorithm has access to a set of historical samples from each arm, and shows that $\Omega(\log T)$ historical samples enable the algorithm to achieve a constant regret. Variants of this problem include allowing the algorithm to receive additional samples in each iteration by paying a certain cost (where the cost is allowed to be zero) [70], or allowing the algorithm to “abstain” in an iteration, so that the sample is observed by the algorithm but not counted towards its regret [67]. In our formulation, the two algorithms receive “auxiliary” data from each other. Contrary to prior work, the treatment algorithm is not free to choose directly what arms are pulled in the auxiliary data which is determined by the control algorithm, and vice versa, though the treatment algorithm indirectly influences the auxiliary data supplied by the control algorithm through data sharing. Such dependencies between the two algorithms do not exist when the historical data is offline.
@@ 1.2.3 Intersection between causal inference and bandit algorithms
Our work lies at the intersection of causal inference, experimentation, and bandit problems. A growing body of prior literature [72, 25, 71, 27, 45, 59] uses bandit algorithms for adaptive data collection to address causal questions that involve evaluating the performance of a policy or an arm. For evaluating policies, one objective is to perform off-policy evaluation, that is, to evaluate the policy’s performance on the true distributions, only given biased data that is collected adaptively. For evaluating arms, the objective is to estimate the mean of each arm or to construct confidence intervals for these means. Prior work examines the bias in the sample mean of each arm due to adaptive sampling [49, 54, 55, 56]. This line of work uses bandit algorithms as a tool to evaluate other objects, whereas our goal is to evaluate bandit algorithms themselves. In other words, we focus on comparing adaptive algorithms rather than fixed policies or arms.
@@ 2.1 Stochastic multi-armed bandit
A bandit algorithm aims to maximize the cumulative reward over a horizon of $T\geq 1$ timesteps. At each timestep $t\in[T]$, the bandit algorithm selects an arm $I_{t}\in[K]$ and receives a stochastic reward $Y_{t}\sim D_{I_{t}}$ from this arm. The arm selection is based on the history of all arm pulls and reward samples in the previous (t1)(t-1) timesteps. Formally, we denote the history at timestep $t$ by $\mathcal{H}_{t}=(\mathcal{H}_{1,t},\ldots,\mathcal{H}_{K,t})$, where $\mathcal{H}_{k,t}$ denotes the set of reward samples collected from arm $k\in[K]$ in the previous (t1)(t-1) timesteps. That is, we define $\mathcal{H}_{k,t}:=\{Y_{i}:I_{i}=k,i\in[t-1]\}$. In words, we consider a class of bandit algorithms that select arms based on the set of reward samples for each arm so far, but are oblivious to the order of the past arm pulls (or reward samples). Many classic bandit algorithms belong to this general class, including the $\epsilon$-greedy algorithm, the Upper Confidence Bound (UCB) algorithm, the EXP3 algorithm, and Thompson sampling. For any algorithm $\mathcal{A}$ in this class, we write the arm selected by this algorithm at timestep $t$ by $I_{t}=\mathcal{A}(\mathcal{H}_{t})$.
@@ 2.2 Bandit algorithms under data sharing
We formulate an A/B experiment comparing two recommendation algorithms as a bandit problem with $2T$ users. We assume that two users arrive at the platform at each timestep $t\in[T]$. These two users are assigned to the treatment or the control algorithm one each uniformly at random. To model data sharing between the treatment and control algorithms, we assume that the two algorithms make recommendations based on past ratings from users in both groups. Formally, consider two bandit algorithms $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ that share history $\mathcal{H}_{t}$ over the entire course of $T$ timesteps. At each timestep $t\in[T]$, each of the two algorithms independently chooses the arm $\mathcal{A}_{1}(\mathcal{H}_{t})$ and $\mathcal{A}_{2}(\mathcal{H}_{t})$ to pull based on history $\mathcal{H}_{t}$, observes a reward sample from the chosen arm, and adds this reward sample to the shared history $\mathcal{H}_{t}$. Thus, two reward samples are added to the shared history at each timestep $t$, and the total number of samples included in history $\mathcal{H}_{t}$ is 2(t1)2(t-1). This procedure of running two algorithms under data sharing is formalized in Algorithm 1.